OpenStudy (anonymous):

Please help me with this question? (my book only has answers for odd numbered exercises) : If y = sin (ax +b) is a general solution to a second order differential equation, find it. So far I got: y' = a cos (ax +b) y'' = - a^2 sin (ax + b) y'' = - a^2 sin (ax + b) + a^2 y = a^2 sin (ax +b) y'' + a^2y = 0 a^2 y = - y'' a^2 = - y'' / y

OpenStudy (somethingawesome):

That looks fine, you can probably leave it as y'' = -a^2 y, or y'' + a^2y = 0.

OpenStudy (anonymous):

No, I can't. I need to find a order 2 DE for which y = sin (ax + b) is the general solution, so far none of the exercises I'm doing have the constants "a" or "b" in them.

OpenStudy (somethingawesome):

I see. Weird. You solved the other problems in the section in a similar way?

OpenStudy (anonymous):

Yes, and in none of them one of the constants "a" or "b" remained in the DE. I could always get something just in terms of f(x) (nought) y'' + f(x) (sub - one) y' + f(x) (sub - two) y = 0

OpenStudy (somethingawesome):

I don't know, this question is frustrating. It does say that "If y = sin (ax +b) is a general solution to a second order differential equation, find it." so maybe it's a trick question and it really isn't a general solution to anything? I'm not sure how to prove it, aside from taking the DE we found and finding what the general solution really is.

OpenStudy (anonymous):

I guess it's ok if the DE has one of the constants in it, it's just that since it didn't happen with any of the others, I thought there might be some way to fin "a" and "b" solely in terms of y, y' or y'', however hard. Well, thanks for your help. :)

OpenStudy (somethingawesome):

Yeah. By the way, I found this: http://www.physicsforums.com/showthread.php?t=227578 They work the same problem and even check that it works for y''+a^2y=0. It's weird, since the constants c_1 and c_2 end up being cos(b) and sin(b) thanks to a trig identity.

OpenStudy (anonymous):

I'll check it.

OpenStudy (anonymous):

Cool! : )